# TextMine: Data, Evaluation Framework and Ontology-guided LLM Pipeline for Humanitarian Mine Action

### v2

@@ A.3 Ontology-Guided for Knowledge Graph Construction
Other line of work focuses on reasoning and validation within already structured knowledge graphs, not extraction from free-form text as with TextMine Liu et al. (2025); Regino and Dos Reis (2025). Their ontological constraints apply post-hoc to existing triples, whereas TextMine uses the ontology during extraction to guide triple formation and reduce hallucinations.
Finally, none of these works address bias-aware evaluation: they rely on deterministic metrics or small-scale manual validation. TextMine introduces a bias-aware LLM-as-Judge evaluation framework that explicitly mitigates positional bias in reference-free scoring—an issue absent from prior ontology-guided pipelines.
@@ A.4 Multi-Faceted Knowledge Graph Evaluation
Appendix B Additional Information and Insights for the HMA Ontology
Figure 7: Number of prompts, entity types, and relation types per the utilized ontologies from IMSMA Core and Empathi.
Appendix C In-Context Learning Prompt Example
@@
Appendix D LLM Judge Prompts
@@
Appendix E Evaluation Reasoning Process

### v3

@@ A.3 Ontology-Guided for Knowledge Graph Construction
Other line of work focuses on reasoning and validation within already structured knowledge graphs, not extraction from free-form text as with TextMineX Liu et al. (2025); Regino and Dos Reis (2025). Their ontological constraints apply post-hoc to existing triples, whereas TextMineX uses the ontology during extraction to guide triple formation and reduce hallucinations.
Finally, none of these works address bias-aware evaluation: they rely on deterministic metrics or small-scale manual validation. TextMineX introduces a bias-aware LLM-as-Judge evaluation framework that explicitly mitigates positional bias in reference-free scoring—an issue absent from prior ontology-guided pipelines.
@@ A.4 Multi-Faceted Knowledge Graph Evaluation
Appendix B In-Context Learning Prompt Example
@@
Appendix C LLM Judge Prompts
@@
Appendix D Evaluation Reasoning Process